### v1

# Data-Adaptive Identification of Subpopulations Vulnerable to Chemical Exposures using Stochastic Interventions

Abstract
In environmental epidemiology, identifying subpopulations vulnerable to chemical exposures and those who may benefit differently from exposure-reducing policies is essential. For instance, sex-specific vulnerabilities, age, and pregnancy are critical factors for policymakers when setting regulatory guidelines. However, current semi-parametric methods for heterogeneous treatment effects are often limited to binary exposures and function as black boxes, lacking clear, interpretable rules for subpopulation-specific policy interventions. This study introduces a novel method using cross-validated targeted minimum loss-based estimation (TMLE) paired with a data-adaptive target parameter strategy to identify subpopulations with the most significant differential impact from simulated policy interventions that reduce exposure. Our approach is assumption-lean, allowing for the integration of machine learning while still yielding valid confidence intervals. We demonstrate the robustness of our methodology through simulations and application to NHANES data. Our analysis of NHANES data for persistent organic pollutants on leukocyte telomere length (LTL) identified age as the maximum effect modifier. Specifically, we found that exposure to 3,3’,4,4’,5-pentachlorobiphenyl (pcnb) consistently had a differential impact on LTL, with a one standard deviation reduction in exposure leading to a more pronounced increase in LTL among younger populations compared to older ones. We offer our method as an open-source software package, EffectXshift, enabling researchers to investigate the effect modification of continuous exposures. The EffectXshift package provides clear and interpretable results, informing targeted public health interventions and policy decisions.
1 Introduction
Identifying vulnerable subpopulations in environmental health research is crucial for understanding how different groups are differentially impacted by chemical exposures and for informing targeted public health interventions. Numerous studies have highlighted that certain populations, such as children, pregnant women, the elderly, and socioeconomically disadvantaged groups, bear a disproportionate burden of environmental pollutants, leading to significant health disparities Balbus and Malina (2009); Ruiz et al. (2018); Zorzetto et al. (2024a).
For example, research consistently shows that African Americans, Latinos, and low-income individuals are more likely to experience higher exposures to endocrine-disrupting chemicals (EDCs) such as polychlorinated biphenyls (PCBs) and bisphenol A (BPA), which are associated with metabolic disorders, including diabetes Ruiz et al. (2018). These groups are disproportionately exposed due to factors such as dietary patterns, the use of consumer products, and residential proximity to pollution sources, particularly air pollution Makri and Stilianakis (2008). Social determinants of health, such as systemic oppression and reduced access to resources, drive these exposure disparities. Additionally, pregnant women and children are particularly susceptible to the toxic effects of EDCs and exposure to metals, especially lead (Pb) Rolfo et al. (2020); Shah-Kulkarni et al. (2020), which can exacerbate health problems such as neurodevelopmental disorders and cardiovascular disease Healey (2009). The per capita healthcare costs associated with health issues related to EDC exposure are significantly higher in these communities, highlighting the urgent need to address this environmental justice crisis Weiss et al. (2023).
Effect modification is a critical concept in epidemiological research, often assessed to understand how different baseline factors influence the relationship between exposure and outcome. Common methods for assessing effect modification include stratified analysis, interaction terms in regression models, regional GIS estimates, interpreting knot points in spline-based or generalized additive models (GAMs), and Bayesian profile regression. For example, Samoli et al. (2011) examined the acute effects of air pollution on pediatric asthma exacerbation by stratifying the analysis by season, sex, and age, using Poisson regression models with interaction terms to investigate the differential effects of pollutants under various conditions, such as during desert dust days. Similarly, Analitis et al. (2014) explored the modification of the effects of heat waves on mortality by air pollution within the EuroHEAT project, employing generalized estimating equation models to assess the joint effects of high temperatures and air pollution on mortality in multiple European cities.
Bayesian methods have shown significant promise in estimating subpopulation effects within environmental health research, particularly with multipollutant exposures. For example, Coker et al. (2018) demonstrated the efficacy of Bayesian profile regression in identifying subpopulations susceptible to different mixtures of pollutants. Zorzetto et al. (2024b) recently introduced Confounder-Dependent Bayesian Mixture Modeling to characterize heterogeneity of causal effects in air pollution studies, leveraging dependent Dirichlet processes to identify population groups with distinct exposure-response relationships for binary exposures.
Recent advancements in heterogeneous effect estimation have also progressed within the domain of semi-parametric statistics. However, many of these meta-learners, such as the T-learner, the S-learner, and the X-learner, often focus on single binary exposures, which are less applicable to environmental epidemiologists dealing with multiple and continuous exposures Künzel et al. (2019). Despite their theoretical strengths, these methods are limited in environmental studies because of their design focus on binary treatments rather than the complex, multidimensional exposures typically encountered. In addition, many act as black boxes for which the analyst chooses which subgroups of individuals to assess treatment effects for, which is often unknown a priori.
The goal of this paper is to introduce a versatile method for causal inference that achieves two key objectives: (i) identifying the exposure-covariate relationships with the maximal differential impact from an intervention, and (ii) estimating the subpopulation intervention effect in the data-adaptively determined subregion using an assumption-lean estimator. This approach integrates machine learning with data-adaptive targeted parameter estimation to provide robust estimates and valid confidence intervals in covariate regions Hub (2016); Zheng and van der Laan (2010). Our methodology is applicable to any type of exposure, including mixed exposures, by employing stochastic shift interventions Muñoz and van der Laan (2012). These interventions alter the exposure distribution and evaluate potential outcomes under this new distribution, extending to mixtures where the exposure or pairs of exposures with the most significant effect modification are identified and estimated.
By simultaneously identifying these relationships and simulating an intervention, our approach necessitates cross-validation to avoid overfitting to our sample data. In this process, the maximal effect modifier is identified in one part of the data, while the effects of policy intervention within the subregion are estimated on a different part of the data. This ensures that our estimates are both reliable and generalizable. If the maximal effect modifier relationship is consistent across data splits, a pooled estimate is calculated leveraging the full data even though the parameter of interest was discovered from the data.
This paper is structured as follows. Section 2 introduces stochastic shift interventions, providing the foundational concepts and notation. In Section 3, we discuss our oracle target parameter, which defines the maximal effect modifier relationship. Section 4 describes the recursive partitioning algorithm designed to estimate this oracle parameter, identifying subregions within the covariates that exhibit the maximal differential impact of a stochastic intervention. Section 5 presents simulations of a mixed exposure scenario, highlighting the identification and accurate estimation of the maximal effect modifier when applying our method. In Section 6, we apply our approach to NHANES data to explore the effects of mixed POP on telomere length, assessing possible effect modification. The code for our method, along with the applications and simulations, is available on GitHub in our open source package called EffectXshift.
2 Stochastic Shift Interventions
Estimating the causal effect of an intervention on a population typically involves defining parameters in a nonparametric structural equation model Pearl (2000), where the treatment or exposure is assigned deterministically. However, in many real-world scenarios, especially in environmental health, exposures are stochastically assigned due to policy interventions that alter the distribution of exposures rather than assigning fixed values. To account for this, we define our new causal parameter for effect modification based on stochastic interventions.
A stochastic shift intervention modifies the distribution of the exposure rather than setting it to a fixed value. This is particularly relevant in environmental health, where interventions aim to reduce the levels of pollutants, say, reduce all people’s exposure by 2 part per million (ppm) rather than setting exposure to a fixed level, which is not feasible realistically. For example, consider a policy that aims to decrease the levels of mixed POPs (such as DDT, polychlorinated biphenyls (PCBs), and dioxins) in the environment to reduce damage to telomere length. Instead of fixing the exposure to a specific level, the policy shifts the distribution of these POPs downward across the population, for example reduce by 1 standard deviation. Therefore, each person still potentially has a different exposure level, but the entire distribution is shifted by some amount, and we then can compare the expected telomere length under this new exposure distribution to the observed outcome under observed distribution. In our proposed method we assume shifts of $\delta$ are known and fixed already, such as for a particular EPA policy it may be known that a fine reduces emissions by Xon average”, or the analyst is interested in reducing an exposure by 1 unit or 1 standard deviation, etc.
Consider an experiment where an exposure variable $A$, a continuous or binary outcome $Y$, and a set of covariates $W$ are measured for $n$ randomly sampled subjects. Let $O=(W,A,Y)$ represent a random variable with distribution $P_{0}$, and $O_{1},\ldots,O_{n}$ represent $n$ i.i.d. observations of $O$. The following NPSEM is assumed to hold:
$W=f_{W}(U_{W});\quad A=f_{A}(W,U_{A});\quad Y=f_{Y}(A,W,U_{Y}),$ (1)
where $U_{W}$, $U_{A}$, and $U_{Y}$ are exogenous random variables and $f$ are nonparametric functions. When $U_{A}\perp U_{Y}$, and either $U_{W}\perp U_{Y}$ or $U_{W}\perp U_{A}$ (randomization assumption) is true, the true distribution $P_{0}$ of $O$ can be factorized as:
$P_{0}(O)=P_{0}(Y|A,W)P_{0}(A|W)P_{0}(W),$ (2)
[…]
$P_{\delta}(g_{0})(A=a|W)=g_{0}(a-\delta|W),$ (3)
for a prespecified $\delta$. This represents a shifted version of the current exposure mechanism, where the shift value is constant across strata defined by the covariates.
The objective is to estimate causal parameters under such stochastic interventions. The causal parameter of interest is defined as:
$P(Y_{P_{\delta}}=y)=\int_{a\in\mathcal{A}}\int_{w\in\mathcal{W}}P(Y_{P_{\delta}}=y|A_{P_{\delta}}=a,W=w)\times P_{\delta}(g)(A=a|W=w)P(W=w),da,dw,$ (4)
Where $A_{P_{\delta}}$ denote the exposure variable under the intervened system (i.e., $A_{P_{\delta}}$ is distributed according to $P_{\delta}(g)$) and where $\mathcal{A}$ and $\mathcal{W}$ are the support of A and W, respectively. This is simply the outcome weighted by the new conditional exposure density under a shift.
From the nonparametric structural equation model (NPSEM) given by equation (1), we know that the conditional probability $P(Y_{P_{\delta}}=y|A_{P_{\delta}}=a,W=w)$ is the same as $P(Y_{a}=y|A_{P_{\delta}}=a,W=w)$. Here, $Y_{a}$ represents the counterfactual outcome when the exposure is set to level $a$ with probability one.
Under the usual randomization assumption, which states that $A\perp Y_{a}|W$, we also have that $A_{P_{\delta}}\perp Y_{a}|W$. This implies that $P(Y_{a}=y|A_{P_{\delta}}=a,W=w)=P(Y_{a}=y|W=w)$ and under the consistency assumption (which means that if $A=a$, then $Y_{a}=Y$), the above quantity is identified by $P(Y=y|A=a,W=w)$ our counterfactual distribution can be written as:
$P(Y_{P_{\delta}}=y)=\int_{a\in\mathcal{A}}\int_{w\in\mathcal{W}}P(Y=y|A=a,W=w)\times P_{\delta}(g)(A=a|W=w)\times P(W=w),da,dw,$ (5)
We define the parameter of interest as a mapping $\Psi:\mathcal{M}\rightarrow\mathbb{R}$ that takes an element in a statistical model $\mathcal{M}$ and maps it into a number in the reals. The true value of the parameter is given by the mapping evaluated at the true distribution $P_{0}\in\mathcal{M}$, and is denoted by $\psi_{0}=\Psi(P_{0})$. Our causal and statistical parameter of interest is then given by:
$E(Y_{P_{\delta}})=\Psi(P)=\int_{A\in\mathcal{A}}\int_{W\in\mathcal{W}}\overline{Q}(A,W)\times P_{\delta}(g)(A|W)\times Q_{W}(W),da,dw,$ (6)
To estimate these quantities in practice, we can use ensemble machine learning methods for $g$ and $\overline{Q}$. These advanced techniques allow us to flexibly and more accurately estimate the necessary components of the model.
2.1 Efficient Estimation and Inference for Stochastic Interventions
To estimate the causal effect of stochastic interventions effectively, we aim to construct an efficient estimator. An efficient estimator is one that achieves the smallest possible asymptotic variance among regular asymptotically linear (RAL) estimators. This ensures that the estimator provides the most precise estimates and the tightest confidence intervals.
An estimator $\hat{\Psi}$ is said to be regular asymptotically linear (RAL) if it satisfies the following property:
$\sqrt{n}(\hat{\Psi}-\Psi)=\sqrt{n}\mathbb{E}_{n}[\phi(O)]+o_{p}(1),$ (7)
where $\phi(O)$ is the influence function. The RAL property implies that the estimator is asymptotically linear, achieving a $\sqrt{n}$ convergence rate, which ensures that the distribution of $\sqrt{n}(\hat{\Psi}-\Psi)$ approaches a normal distribution with mean zero and variance equal to the expectation of the squared influence function. Efficient estimators are those RAL estimators that achieve the smallest possible asymptotic variance among all RAL estimators.
where $\phi(O))$ is the influence function, and $\mathbb{E}_{n}$ denotes the empirical mean. By the central limit theorem, RAL estimators achieve the asymptotic variance bound $\mathbb{E}[\phi^{2}(W)]$. This implies that as the sample size increases, the distribution of $\sqrt{n}(\hat{\Psi}-\Psi)$ approaches a normal distribution with mean zero and variance $\mathbb{E}[\phi^{2}(W)]$, facilitating valid statistical inference.
In the context of our stochastic shift intervention, we seek an estimator $\hat{\Psi}$ for the causal parameter $\Psi$. The discrepancy between the estimator and the true parameter, $\Psi-\hat{\Psi}$, can be characterized using an influence function. The influence function not only quantifies the sensitivity of the estimator to individual observations but also provides a way to derive the asymptotic variance of the estimator, which is crucial for constructing valid confidence intervals.
For our specific stochastic intervention, the influence function $\phi(O)$ is given by:
$\phi(W)=\frac{g(A-\delta|W)}{g(A|W)}\left(Y-\overline{Q}(A,W)\right)+\overline{Q}(A+\delta,W)-\Psi(P).$ (8)
This function represents the deviation of an observation $O=(W,A,Y)$ from the parameter $\Psi(P)$, accounting for the shift $\delta$ in the exposure distribution. Intuitively, the influence function describes how the estimator $\hat{\Psi}$ behaves as we incorporate new observations into our sample. Specifically, the first term $\frac{g(A-\delta|W)}{g(A|W)}\left(Y-\overline{Q}(A,W)\right)$ captures the adjustment for the shift in exposure, weighted by the relative likelihood of the shifted versus original exposure. The second term, $\overline{Q}(A+\delta,W)$, adjusts the expected outcome under the shifted exposure, and the final term $\Psi(P)$ centers the influence function around the true parameter value.
Various estimators can be employed for efficient estimation of our stochastic shift target parameter, including inverse probability weighting (IPW), augmented inverse probability weighting (AIPW), and targeted learning. Among these, we choose targeted learning to estimate our shift parameter due to its robustness in finite samples Li et al. (2022).
The IPW estimator uses the estimated propensity score:
$H(A,W)=\frac{g(A|W)}{g(A-\delta|W)}$ (9)
to weigh observations, allowing for the estimation of causal effects. However, it is inefficient and relies heavily on the correct specification of the propensity score model, which in this case is a density estimator, a particularly difficult estimator where there are not many flexible and reliable methods. As such, this estimator may be unreliable given our need to reliably estimate the conditional density.
The AIPW estimator combines IPW with outcome regression, providing double robustness. It remains consistent if either the propensity score model or the outcome model is correctly specified. However, this estimator is not a plug-in estimator, and therefore does not always respect the bounds of the outcome thereby resulting in potentially more variability.
Targeted Maximum Likelihood Estimation (TMLE) offers a robust and efficient approach by iteratively updating initial estimates through a least favorable submodel. This method adjusts for potential biases in the initial estimates using the clever covariate, ultimately providing more accurate estimates in lower sample sizes. This approach results in a plug-in esitmator which respects the bounds of the data, resulting in lower variability in finite samples.
Targeted learning begins with initial estimates of the outcome regression ($Q$) and the propensity score ($g$), which can be obtained through flexible machine learning techniques. These initial estimates are then updated using a clever covariate, which is constructed to capture the variation related to the shift parameter, $\delta$. The refinement process involves fitting a parametric fluctuation model that minimizes a specified loss function. This iterative process continues until convergence, resulting in a final estimate that is both efficient and less biased. This debiased esitmate is then plugged into the influence function where we can derive Wald-style confidence intervals.
3 Causal Subpopulation Intervention Effects
Our primary objective in this paper is not only to estimate the overall effects of environmental exposures but also to discern how these effects differ across various sub-populations. Within the framework of stochastic shift interventions, our focus is on identifying specific types of individuals from baseline characteristics that change the impact of policy interventions, such as reducing exposure to pollutants. Unlike the previously described stochastic shift parameter, which averages effects across the entire population, we now aim to define our causal parameter for particular subgroups. This allows us to pinpoint the subpopulations where the intervention yields the maximum differential effect, thereby providing more targeted and effective public health strategies.
3.1 Oracle Parameter for Maximal Difference in Expected Outcomes in Subregion
The oracle parameter, denoted as $\psi^{*}$, represents the maximal differential impact of an intervention in a specific subregion of the covariate space. It is defined as follows:
$\psi^{*}=\underset{A_{i}\in\boldsymbol{A},\,\mathcal{V}\subseteq W}{\text{max}}$$\left\{\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}\right]\right.$ (10)
$\left.-\mathbb{E}\left[Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y\mid W\in\mathcal{V}^{c}\right]\right\},$ (11)
where $A_{i}$ represents an individual exposure within the mixed exposure set $\boldsymbol{A}$. $\mathcal{V}$ and $\mathcal{V}^{c}$ denote subregions of the covariate space $W$. $Y_{A_{\delta_{i}}}$ are the potential outcomes under the specified conditions. The goal is to identify the combination of $A_{i}$ and $\mathcal{V}$ that maximizes the difference in expected outcomes, indicating a modifying effect of the covariate on exposure.
To motivate why we care about the contrast between $W\in\mathcal{V}$ and $W\notin\mathcal{V}$ instead of just looking at the shift effect in the $\mathcal{V}$ region, consider the following. Let $Z$ denote the difference in potential outcomes:
$Z=Y_{A_{i}-\delta_{i},\mathbf{A}_{-i}}-Y$
Then, our contrast can be expressed as:
$\mathbb{E}[Z\mid W\in\mathcal{V}]-\mathbb{E}[Z\mid W\notin\mathcal{V}]$
Denote $W\in\mathcal{V}$ as $W=1$ and $W\notin\mathcal{V}$ as $W=0$. By the law of total expectation, the overall average effect $m$ can be expressed as:
$m=\mathbb{E}[Z]=\mathbb{E}[Z\mid W=1]P(W=1)+\mathbb{E}[Z\mid W=0]P(W=0)$
Let $m_{1}=\mathbb{E}[Z\mid W=1]$ and $m_{0}=\mathbb{E}[Z\mid W=0]$. The contrast we care about can be rewritten as:
$m_{1}-m_{0}=\frac{m_{1}-m}{P(W=0)}$
The overall average $m$ is fixed. Therefore, comparing to just $m_{1}$, our contrast will pick regions where $P(W=0)$ is small and $m_{1}$ is large, as opposed to just trying to maximize $m_{1}$. This approach ensures that we do not end up selecting a region that is too small or insignificant in size, which would typically result in a single point, but rather find a meaningful subregion that exhibits a significant differential impact.
In general, we are interested in the causal exposure-covariate region combination where the effect of the intervention, compared to the complementary covariate space, is maximally different, controlling for other exposures. From now on, we refer to these regions as $V$ and $V^{c}$.
This method allows us to identify and focus on subpopulations where the intervention has the most substantial and relevant impact, thereby informing targeted public health strategies and interventions.
3.2 Statistical Subpopulation Intervention Effects
Here, we explain how to do inference when the data-adaptive parts are fixed, that is, assuming $V,V^{c}$ are fixed. With the causal and statistical parameter determined, we then explain how the regions are generated. Given the same assumptions as stochastic shift interventions, our statistical parameter $\psi$ estimates our causal parameter and is defined as:
$\psi$ $=\underset{A_{i}\in\boldsymbol{A},\,V\subseteq W}{\text{arg max}}\Bigg{(}\mathbb{E}\left[\mathbb{E}\left[Y\mid A_{i}-\delta_{i},\mathbf{A}_{-i},W\in V\right]-\mathbb{E}[Y\mid W\in V]\right]$ (12)
$\quad-\mathbb{E}\left[\mathbb{E}\left[Y\mid A_{i}-\delta_{i},\mathbf{A}_{-i},W\in V^{c}\right]-\mathbb{E}[Y\mid W\in V^{c}]\right]\Bigg{)},$ (13)
This parameter of effect modification, $\Psi_{V}(P)$, is defined as the expected outcome under a stochastic intervention on exposure $A_{i}$, specifically within a subregion $V$ of the covariate space. It integrates over all possible values of $W$ within the subset $V$, capturing the differential impact of $A_{i}$ on the outcome across covariate subgroups.
Operationally, this requires our algorithm to search the exposure-covariate space and, for each exposure, measure the expected outcome change under a shift in regions of the covariates. The algorithm then selects the region and its complementary region where the difference in stochastic intervention impact is maximally different.
The above statistical parameter matches the oracle parameter of interest. However, we may want to include the variance of estimates within each region in order to compare the regions and search for the region difference with the largest statistically significant difference, such as the effect weighed by standard error. To do this, we calculate the influence function in each region, defined below:
$\phi(W,V)=\frac{g(A-\delta|W\in V)}{g(A|W\in V)}\left(Y-\overline{Q}(A,W\in V)\right)+\overline{Q}(A+\delta(W),W\in V)-\Psi(P,V),$ (14)
$\phi(W,V^{c})=\frac{g(A-\delta|W\in V^{c})}{g(A|W\in V^{c})}\left(Y-\overline{Q}(A,W\in V^{c})\right)+\overline{Q}(A+\delta(W),W\in V^{c})-\Psi(P,V^{c}).$ (15)
These influence functions are calculated within the regions $V\in\mathcal{W}$ and $V^{c}\in\mathcal{W}$, allowing us to estimate the effect modification parameter based on a statistical test of differences, like a simple t-statistic.
Although the oracle parameter is defined in terms of population-level effects where everything is perfectly significant, in practical applications, considering statistical significance can provide more robust measures of region selecting in finite samples. Although this approach does not correspond directly to the oracle parameter, it is meaningful to the analyst as it ensures that the identified regions exhibit both substantial and statistically significant differences.
In the method implementation, the package EffectXshift, users can choose whether to find regions with maximum point estimate difference or based on a t-statistic, moving forward we use the t-statistic approach in simulations and applications.
3.3 Identifying Subpopulations with Maximal Effect Difference
Estimation of the subpopulation intervention effect involves a two-stage procedure. In the first stage we obtain efficient estimation of the individual intervention effects (the expected outcome change for each person, which we denote as IIE) and influence function estimates (IFE) for each exposure and the second stage where we use these estimates in a greedy partitioning algorithm to find the exposure-covariate region with the maximum effect modification.
First Stage: In the first stage, we obtain efficient estimates of the IIE for each exposure $A_{i}$ in $\mathbf{A}$. This is achieved by applying a shift intervention $\delta_{i}$ and debiasing our predictions using TMLE and calculating the expected difference between the outcome under an exposure change compared to no change. Concurrently, we estimate the IFE estimates for each exposure, which quantifies the impact of an individual on our estimator for a shift in the respective exposure.
Specifically, we iterate through the exposures to construct a matrix of IIEs, with each column corresponding to a different exposure. In addition, we create a matrix of IFEs. The first stage consists simply of estimating a stochastic shift intervention for each exposure in the mixture. This is done by 1. Fitting a Super Learner (ensemble of machine learning algorithms that uses cross-validation to select the best estimator or create a linear combination of estimator predictions) for $E[Y|\boldsymbol{A},W]$ which is used to create initial predictions for $\mathbb{E}[Y|A_{i}-\delta,W]$. 2. We then fit a density estimator (or classification estimator - described later) for $g(A_{i}|W)$ and get predictions for $g(A_{i}|W)$ and $g(A_{i}-\delta_{i}|W)$ (or the ratio directly) 3. We then debias our initial predictions for $\mathbb{E}[Y|A_{i}-\delta,W]$ using a least favorable submodel (TMLE) that fits a quasi-binomial regression, regressing the observed outcome (scaled to 0 and 1) onto the clever covariate estimates under no shift offset by the initial predictions from the Super Learner. Using this model, we then input the initial predictions given a shift and density given a shift and get updated predictions through the model which are debiased towards our shift parameter of interest. We then plug these estimates in our influence function to get the ICE for each observation.
We loop through each exposure in the mixture and get debiased estimates of the effect of a shift $\delta_{i}$ in each exposure $A_{i}$, constructing a matrix of individual intervention effects. We also construct a matrix for the influence curve estimates for each intervention in each exposure.
Second Stage: In the second stage, we loop through the IIE and IC estimates for each exposure and regress these onto the covariates using a greedy partitioning algorithm. This algorithm searches the covariate space and finds the region and it’s complement which has the maximum effect difference based on a t-statistic, comparing the averages from the IIE and standard error estimates from the IC in each covariate region and splitting based on decreasing p-value. The detailed pseudocode in Algorithm 1 describes the partitioning approach, using the t-statistic to find the exposure-covariate region that has the maximum difference in subpopulation intervention effect.
Algorithm 1 Recursive Covariate Partitioning Based on T-Statistic
1:Input: Data $D$, covariates $W$, individual intervention effect $IIE$, influence curve estimates $ICE$, max depth $d_{\text{max}}$, minimum observations $min\_obs$
2:Output: Rule that demarcates the covariate region with the expected outcome difference
3:procedure T-Part($D$, $W$, $IIE$, $ICE$, depth=0)
4:     if depth $=d_{\text{max}}$ then
5:         return $best\_split$triangleright Stop recursion
6:     end if
7:     Init $best\_split\leftarrow\text{Null}$, $p_{\text{best}}\leftarrow 1$
8:     for $w\in W$ do
9:         for $s$ in unique values of $w$ do
10:              $\mathcal{W}^{*}\leftarrow(D[w]\leq s)$
11:              if sum($\mathcal{W}^{*}$) ¡ $min\_obs$ or sum($\neg\mathcal{W}^{*}$) ¡ $min\_obs$ then
12:                  continue
13:              end if
14:              Split $D$ into $D_{\text{left}}$ and $D_{\text{right}}$ based on $\mathcal{W}^{*}$
15:              Calculate mean IIE for $D_{\text{left}}$ and $D_{\text{right}}$
16:              Calculate variance of ICE for $D_{\text{left}}$ and $D_{\text{right}}$
17:              Calculate t-statistic using:
[…]
18:              Calculate p-value from t-stat
19:              if p-value $<p_{\text{best}}$ then
20:                  $p_{\text{best}}\leftarrow p\_value$
21:                  $best\_split\leftarrow\{w,s,p_{\text{best}}\}$
22:              end if
23:         end for
24:     end for
25:     if $best\_split\neq\text{Null}$ then
26:         Split $D$ into $D_{\text{left}}$ and $D_{\text{right}}$ based on $best\_split$
27:         Recursively apply T-Part($D_{\text{left}},W,IIE,ICE,depth+1$)
28:         Recursively apply T-Part($D_{\text{right}},W,IIE,ICE,depth+1$)
29:     else
30:         return No more significant splits found $\triangleright$ Stop recursion
31:     end if
32:end procedure
33:Main Loop:
34:for $IIE_{i}\in 1:\text{ncol}(IIE)$ do
35:     Initiate T-Part($D$, $W$, $IIE_{i}$, $ICE_{i}$, 0)
36:end for
37:Select the top max effect modifier-covariate relationship: extract region rule
38:Return $max\_mod\_rule$
3.4 Cross-Validation
As discussed, sample splitting is required to avoid bias incurred by discovering subregions of the covariate space and estimating the effect of subpopulation intervention using the same data. For this we use K-fold cross-validation in which the sample data are split into K-folds of roughly equal size. For example, consider $K$ = 10, and define $k=1$. The $K-1/K$ portion of the sample is used to define the effect modification parameter of interest (here 90%) and the $1/K$ (10%) is used to estimate the parameter, the subpopulation intervention effect. Let $T_{n,k}\subset\{1,\ldots,n\}$ denote the observations indices in the k-th training sample and $V_{n,k}$ denote the observations indices in the k-th validation sample for $k=1,\ldots,K$. The following paragraphs describe how through cross-validation the $T_{n,k}$ (training) portion of the data is used to identify the maximal exposure-effect modifier sets in the mixture and $V_{n,k}$ (validation) is used to estimate our subpopulation stochastic shift intervention parameter.
Section 3.3 described our approach to discovering the covariate subregion with the maximum difference in the intervention effect, this is done using the $T_{n,k}$ training data. More specifically, we train our Q and g estimators using this training data and get targeted estimates for the IIE and ICE. These are then regressed onto our t-statistic tree algorithm to find the region in the covariate space with the largest significant difference in average change due to intervention. The result of this process is a rule such as Sex $=$ 1 or Age $>$ 52 and BMI $>$ 20, etc. which is the region that has the maximum effect difference compared to the complement to that region found in the training estimates.
Now, treating this rule as fixed, we get targeted predictions for our IIE and ICE in the validation sample ($V_{n,k}$) from the Q, g, and targeting models trained in the training fold. We then subset the validation data based on the maximum effect modifying rule found in training and simply take the average of the validation IIE to get point estimates and calculate standard errors from the ICE in the subregions indicated by the rule.
This process is done for each fold of the data and therefore may result in different regions found to be the maximum effect modifying region set.
3.5 Pooled TMLE
Upon completion of the cross-estimation procedure, a pooled TMLE update provides a summary measure of the oracle maximum effect modification parameter across k-folds. In each fold, we have estimates for Q and g for the respective regions found to be the maximum effect modifying region and it’s complement (again, these may vary). For the pooled TMLE update we stack the estimates for these nuisance parameters from the estimation samples and run a pooled TMLE update on the cumulative initial estimates. The resulting average is then used for parameter estimation. The pooled estimate across the folds can be succinctly represented as:
$\theta=\frac{1}{K}\sum_{k}\Psi_{F_{T_{n,k}}}\left(V_{n,k}\right)(P_{0})$ (17)
Where $\Psi_{F_{T_{n,k}}}$ identifies optimal exposure-covariate modifier set and subsequently formulates a plug-in estimator using training data $T_{n,k}$. $V_{n,k}$ refers to the estimation sample of the data for deriving our estimates. Essentially, this provides an average estimate for the regions $v$ and $v^{c}$ found to have the maximum intervention effect difference across the folds. If $v$ and $v^{c}$ correspond to the same region across all the folds, such as consistently finding age $<22$ for $v$ and age >=>=> = 22 for $v^{c}$ then these pooled estimates are interpretable and provide more power in the regional estimates as they leverage the full data, resulting in tighter confidence intervals.
3.6 Pooled TMLE for Marginal Exposure Effects
In the case where exposure is a mixture, it is often of interest to the analyst to understand the importance for each exposure relative to the others. Because our approach requires estimation of a shift on each exposure and derivation of the ICE in order to search for the exposure-covariate region with maximum effect modification, we easily can provide CV-TMLE estimates of stochastic shift. That is, for each exposure we simply take our $Q$ and $g$ estimators, trained on the training data, and get the predicitons for the validation data. We then stack our initial estimates $Q$ and $g$, do a pooled TMLE on the full data, and update our counterfactuals of the change in the outcome due to a change in each exposure. This provides estimates for the impact of an exposure change for each exposure with valid inference and can be used to assess for exposure importance.
4 Density Estimation of the Exposures
Estimating the density of each exposure is a crucial component of our methodology to evaluate the effects of the intervention both on average across the population and in subpopulations. We provide two approaches to density estimation, direct density estimation, and a reparameterization based on classification, depending on user input.
4.1 Direct Density Estimation
Super Learner Algorithm: Super Learner operates by training a variety of base learners (individual algorithms) on the data and then combining their predictions using a meta-learner. The meta-learner, in this case, is a non-negative least squares (NNLS) algorithm, which assigns weights to each base learner based on their performance. This approach leverages the strengths of different algorithms to produce a more accurate and robust predictive model.
Homoscedastic Errors (HOSE) Learner: The HOSE learner assumes that the variance of the errors is constant (homoscedasticity) across the range of the predicted values. It employs a semi-parametric approach to density estimation, where the mean of the distribution is estimated using a Super Learner composed of multiple base learners.
Heteroscedastic Errors (HESE) Learner: The HESE learner, in contrast, allows for varying error variances (heteroscedasticity) across different levels of the predicted values. It employs a two-step semi-parametric approach. The mean of the distribution is estimated using a similar Super Learner as in HOSE. An additional Super Learner is used to model the variance of the errors, allowing the error variance to change depending on the value of the predictors.
Although direct estimation is straightforward in principle, it can be challenging in practice due to the scarcity of effective data-adaptive estimators for high-dimensional conditional densities in the machine learning and statistical literature. As such, we also provide another approach.
4.2 Reparameterization of Density Estimation via Classification
As an alternative to direct density estimation, we propose using a novel approach that redefines the problem of estimating the density ratio $H(A,W)$ as a classification problem in an augmented data set [CITE]. This method takes advantage of the fact that classification algorithms are well-developed and widely available in the machine learning literature compared to density estimators.
To apply this approach, we construct an augmented dataset consisting of duplicated observations. For each original observation, we create a duplicate with the observed exposure $A_{t}$ and another with the exposure under the intervention $A^{d}_{t}$. We introduce an indicator variable $\xi$, which is set to 1 if the duplicated observation corresponds to the exposure under intervention, and 0 otherwise. The augmented data set can be represented as:
$\{(W_{\xi,i,t},A_{\xi,i,t},\xi_{\xi,i}):\xi=0,1;i=1,\ldots,n\},$
where $\xi_{\xi,i}=\xi$ indexes the duplicates, $W_{\xi,i,t}=W_{i,t}$ is the history variable, and $A_{\xi,i,t}=\xi\cdot A^{d}_{i,t}+(1-\xi)\cdot A_{i,t}$ is the natural exposure level if $\xi=0$, and the intervened exposure level if $\xi=1$.
We denote the probability distribution of $(W_{t},A_{t},\xi)$ in the augmented dataset by $P_{\xi}$, with corresponding density $p_{\xi}$. We define the parameter $u_{\xi}$ of $P_{\xi}$ as:
$u_{\xi}(a_{t},w_{t})=P_{\xi}(\xi=1|A_{t}=a_{t},W_{t}=w_{t}).$
The density ratio $H(A,W)$ can be expressed in terms of $u_{\xi}$ as follows:
$H_{t}(a_{t},w_{t})=\frac{p_{\xi}(a_{t},w_{t}|\xi=1)}{p_{\xi}(a_{t},w_{t}|\xi=0)}=\frac{P_{\xi}(\xi=1|A_{t}=a_{t},W_{t}=w_{t})}{P_{\xi}(\xi=0|A_{t}=a_{t},W_{t}=w_{t})}=\frac{u_{\xi}(a_{t},w_{t})}{1-u_{\xi}(a_{t},w_{t})},$
where the first equality follows from the definition of $r_{t}$ and the definition of conditional density, the second from Bayes’ rule and the observation that $P_{\xi}(\xi=1)=P_{\xi}(\xi=0)=\frac{1}{2}$, and the third from the definition of $u_{\xi}$.
By estimating $u_{\xi}$ in the augmented dataset using any classification method available in the machine learning and statistical learning literature (e.g., Super Learning), we can indirectly estimate the density ratio. This classification-based approach ensures that the estimator retains desirable properties such as asymptotic normality, provided that appropriate cross-validation and cross-fitting techniques are employed. Furthermore, this reclassification approach is substantially faster computationally which is necessary, especially in our proposed methodology, where the conditional density is needed for each exposure and higher folds in cross-validation provide more consistent results. [CITE]
In our implementation, users can choose between direct density estimation and the reparameterization approach via a user-defined parameter.
5 Summary of Effect Modification Algorithm
Our method of identifying subpopulations differentially impacted by mixed exposures involves several key steps: 1. Identification of Subpopulations: We aim to identify subpopulations that are differentially impacted by mixed exposures. This is achieved by establishing an oracle target parameter that defines the maximal effect modifier relationship. This oracle parameter serves as the benchmark for identifying the subpopulations with the most significant differential impacts from an intervention. 2. Estimation of Oracle Parameter: Once the oracle target parameter is defined, we estimate it by finding the subregions within the covariate space that best approximate our oracle. This involves developing a t-statistic tree, which is a recursive partitioning algorithm that finds the significantly different subregions based on targeted counterfactuals of a shift intervention and influence curve estimates for each exposure. 3. Recursive Partitioning Algorithm: Our recursive partitioning algorithm, guided by t-statistics, identifies the subregions where the impact of the intervention is maximally different. By comparing the averages from the individual intervention effects (IIE) and standard error estimates from the influence curves (IC) in each covariate region, the algorithm splits based on decreasing p-values. 4. Honest Estimation through Cross-Validation: To ensure honest estimation and avoid overfitting, we employ cross-estimation. We use training folds to generate the regions that match our oracle target parameter and estimation samples to estimate the impact of intervention in those regions. This cross-validation process ensures that our results are reliable and generalizable. 5. Efficient Estimation with TMLE: We use TMLE to obtain efficient estimates of the individual intervention effects for each exposure. This method provides robust estimates and valid confidence intervals, leveraging machine learning to flexibly and accurately estimate the necessary components of the model. 6. Pooled Estimates for Robust Inference: After completing the cross-validation procedure, we perform a pooled TMLE update to provide a summary measure of the oracle maximum effect modification parameter across folds. This pooled estimate combines the information from all folds for the regions $v$ and $v^{c}$.
6 Simulation Study
6.1 Data Generation with Binary Modifiers
To evaluate the performance of our proposed method, we conducted a simulation study using synthetic data that mirrors a realistic environmental epidemiology scenario. We generated a large population with 10000 observations to estimate ground truth of the subpopulation intervention effect. Three binary confounders, $W_{1}$, $W_{2}$, and $W_{3}$, were independently generated from Bernoulli distributions with a success probability of 0.5.
The exposures ($A_{1},A_{2},A_{3}$) were generated as follows:
$A_{1}$ $\sim\mathcal{N}(0.5W_{3}+0.3W_{2}+0.4W_{1},1)$
$A_{2}$ $\sim\mathcal{N}(0.3W_{2}+0.3W_{1},1)$
$A_{3}$ $\sim\mathcal{N}(0.2W_{1},1)$
The outcome $Y$ was then simulated based on the exposures and covariates, incorporating effectm modification effects:
$Y=2+1\cdot A_{1}+0.5\cdot A_{2}+0.2\cdot A_{3}+0.5\cdot W_{3}-0.3\cdot W_{2}+0.4\cdot W_{1}+2\cdot A_{1}\cdot W_{3}+\epsilon$
where $\epsilon\sim\mathcal{N}(0,1)$ represents random noise.
6.2 Data Generation with Continuous Modifier
To make detection of the modifying region more difficult and arguably more realistic compared to the binary case, we also create a DGP with a continuous covariate where the impact on the outcome jumps at a threshold. For this, we change $W_{3}$ to a continuous $\mathcal{N}(40,10)$ and create a binary indicator for when $W_{3}>55$ and use this indicator in place of $W_{3}$ in the above outcome generation equation, but provide the EffectXshift algorithm the continuous covariate to determine whether it detects the correct threshold that leads to modification.
6.3 Performance Metrics
To establish the ground truth for the differential impacts of exposures, we applied a shift intervention to reduce $A_{1}$ by a fixed amount ($\delta=-0.5$) and recalculated the outcome $Y$ to obtain $Y_{\delta}$. The difference between the shifted and unshifted outcomes ($\Psi=Y_{\delta}-Y$) was aggregated to estimate the true average effect within each level of the covariate $W_{3}$, where the true effect modification was created. Here the true average in level 0 for $W_{3}$ is region $V$ and level 1 is $V^{c}$.
We assessed the bias, variance, mean squared error (MSE), and coverage of our method across different sample sizes ($n=300,500,1000,2000,5000$). For each sample size, we ran the simulation 100 times to evaluate the stability and reliability of our estimator. Bias: The average difference between the estimated and true effect. Variance: The variability of the estimator across simulations. MSE: Sum of the squared bias and variance, providing a comprehensive measure of estimator accuracy. Coverage: The proportion of times the true effect lies within the estimated 95% confidence interval.
These metrics are essential for ensuring the $\sqrt{n}$-consistent behavior of our estimator, meaning that as the sample size increases, the estimator converges to the true parameter value at a rate proportional to the square root of the sample size. This property is crucial for an efficient estimator, as it guarantees precise and reliable estimates in large samples.
For our estimator to be asympotically unbaised we need to accurately detect the correct region with increasing sample size. As such we report metrics for Accuracy, Precision, Recall and F1 statistics as well.
6.4 Simulation Process
The simulation process involved the following steps: 1. Data Generation: For each sample size, we generated datasets of the above data. 2. Estimator Application: We applied our method to each dataset, to estimate the intervention effects within discovered subregions using 10-fold CV and default machine learning algorithms in the Super Learners (xgboost, random forest, glm) for the outcome regression and density estimation. 3. Metric Calculation: For each sample size and iteration, we calculated the bias, variance, MSE, and coverage of the estimated effects. 4. Result Aggregation: The results were aggregated to evaluate the overall performance of our method across different sample sizes.
7 Simulation Results
7.1 Identification of Maximum Effect Modifying Subregion
For the binary effect modifier, for all iterations and sample sizes, we consistently identified the correct modifying variable, $W_{3}$ for the correct region, therefore accuracy, precision, recall and f1 were all 1.0.
The performance of our method in identifying the true subpopulation in continuous data improves as sample size increases. At a sample size of 300, the accuracy is 59.1%, precision is 18.9%, recall is 100%, and the F1 score is 29.9%. The high recall indicates perfect sensitivity, but the low precision suggests a high number of false positives. With a sample size of 500, accuracy improves to 94.2%, precision to 56.5%, recall remains at 100%, and the F1 score rises to 71.2%. The precision increase indicates fewer false positives. At 1000 samples, accuracy is 96.9%, precision is 81.8%, recall is 90.9%, and the F1 score is 82.9%. The balance between precision and recall reflects improved detection. For 2000 samples, accuracy is 99.7%, precision is 96.4%, recall is 99.3%, and the F1 score is 97.8%, showing minimal false positives and false negatives. At 5000 samples, the method nearly perfects performance with an accuracy of 99.9%, precision of 99.6%, recall of 99.1%, and an F1 score of 99.3%. Figure 1 shows these results.
Figure 1: Subpopulation Detection Metrics for Continuous Modifier
7.2 Binary Effect Modifier Simulation Results
The results are summarized below based on bias, variance, mean squared error (MSE), and coverage probability.
  • Bias and Variance: The bias and variance generally decreased as the sample size increased. For smaller samples (e.g., $n=300$), the bias was more pronounced in the measure in the $v^{c}$ region compared to the $v$ region, suggesting potential overfitting when sample sizes were limited. However, as the sample size increased, the bias significantly reduced, particularly evident in the $v^{c}$ region at $n=5000$ where the bias was minimal ($-0.007$).
  • Mean Squared Error (MSE): The MSE consistently decreased as the sample size increased, reflecting improvements in estimator accuracy. For instance, the MSE in the $v$ region decreased from 0.017 at $n=300$ to 0.0001 at $n=5000$. Similarly, in the $v^{c}$ region, the MSE decreased from 0.010 at $n=300$ to 0.0001 at $n=5000$.
  • Coverage Probability: The coverage probability was stablazlied to 0.95 for the $v$ region across sample sizes but $v^{c}$ region maintained a coverage probability of 1, indicating overly conservative estimates which may be due to limited iterations or our DGP.
7.3 Continuous Effect Modifier Simulation Results
The results of the continuous effect modifier simulation are summarized below based on bias, variance, mean squared error (MSE), and coverage probability.
  • Bias and Variance: As the sample size increased, both bias and variance generally decreased. For the $v$ region, the bias reduced from 0.155 at $n=300$ to a minimal 0.001 at $n=5000$. Similarly, in the $v^{c}$ region, the bias decreased from 1.025 at $n=300$ to 0.112 at $n=5000$, indicating significant improvement in accuracy with larger sample sizes.
  • Mean Squared Error (MSE): The MSE consistently decreased as the sample size increased, reflecting improvements in estimator accuracy. For the $v$ region, the MSE dropped from 0.026 at $n=300$ to 0.0007 at $n=5000$. In the $v^{c}$ region, the MSE decreased from 1.087 at $n=300$ to 0.014 at $n=5000$, demonstrating a substantial reduction in prediction error with larger samples.
  • Coverage Probability: The coverage probability showed consistent improvement with increasing sample size. For the $v$ region, coverage increased from 0.80 at $n=300$ to 0.95 at $n=5000$, indicating better confidence interval reliability. In the $v^{c}$ region, coverage improved from 0.72 at $n=300$ to 0.95 at $n=5000$, suggesting that the true parameter value was increasingly captured within the calculated confidence intervals as sample size grew. Variability in coverage may be due to limited iterations in our simulation.
Overall, the results suggest that our estimator’s mean squared error decreases at a rate proportional to $\sqrt{n}$, aligning with theoretical expectations.
Figure 2: Simulation Results for Binary Effect Modifier
Figure 3: Simulation Results for Continuous Effect Modifier
7.4 EffectXshift Applied to NHANES Dataset
The 2001-2002 National Health and Nutrition Examination Survey (NHANES) cycle provides a robust data set for our analysis, widely recognized for its credibility in the public health domain. This data set includes interviews with 11,039 individuals, of whom 4,260 provided blood samples and consented to DNA analysis. For our analysis, we focused on a subset of 1,007 participants, ensuring complete exposure data. This is comparable to the subset used by Mitro et al. Mitro et al. (2016), who investigated the association between persistent exposure to organic pollutants (POPs) binding to the aryl hydrocarbon receptor (AhR) and leukocyte telomere length (LTL). This data set was also used by Gibson et a., Gibson et al. (2019) who applied various mixture methods to this data. For these reasons, this data set works nicely for comparison of the marginal variable importance effects given our stochastic shift interventions and provides evaluation of effect modification which has otherwise not been assessed for in these methods.
Our study examined the effect of 18 congeners, including 8 non-dioxin-like PCBs, 2 non-ortho PCBs, 1 mono-ortho PCB, 4 dioxins, and 4 furans. All congeners were adjusted for lipid content in serum using an enzymatic summation method. Leukocyte telomere length was measured using quantitative polymerase chain reaction (qPCR) to determine the T/S ratio, which compares telomere length to a standardized reference DNA.
Our analysis considered several covariates: age, sex, race/ethnicity, education level, BMI, serum cotinine levels, and blood cell distribution and count. Categories for race/ethnicity, education, and BMI were consistent with previous studies Mitro et al. (2016); Gibson et al. (2019). Given evidence that telomere length can be influenced by factors such as smoking and age Valdes et al. (2005); Hoefnagel et al. (2016), our approach aimed to identify baseline covariate regions and specific exposures that exhibit the maximum differential impact due to a shift intervention. All covariate were included as potential effect modifiers.
We utilized our EffectXshift package to perform this analysis. Using 10-fold cross-validation (CV), we explored the counterfactual changes in telomere length for a reduction of each exposure by 1 standard deviation. The standard deviation values for each exposure are as follows: PCB74 Lipid Adj (ng/g): -13589.77, 1,2,3,6,7,8-hxcdd Lipid Adj (pg/g): -40.36301, 1,2,3,4,6,7,8-hpcdd Lipid Adj (pg/g): -55.36449, 2,3,4,7,8-pncdf Lipid Adj (pg/g): -5.758755, 1,2,3,4,6,7,8-hxcdf Lipid Adj (pg/g): -10.67117, and 3,3’,4,4’,5-pcnb Lipid Adj (pg/g): -52.7907.
7.5 NHANES Data Results
We evaluated the impact of 16 persistent organic pollutants (POPs) on leukocyte telomere length (LTL). Our primary objective was to identify the baseline covariate region and the exposure with the maximum differential impact due to a reduction in exposure.
7.6 Identification of Maximum Effect Modifiers
The shift interventions were applied to a subset of uncorrelated exposures: PCB74, 1,2,3,6,7,8-hxcdd, 1,2,3,4,6,7,8-hpcdd, 2,3,4,7,8-pncdf Lipid Adj, 1,2,3,4,6,7,8-hxcdf, and 3,3’,4,4’,5-pcnb. Age was consistently identified as the maximum effect modifier across all analysis folds. The exposure 3,3’,4,4’,5-pcnb (LBXPCBLA) was found in 100% of the folds, indicating a robust association. Table 1 shows the K-fold results. The change in outcome due to a 1 standard deviation reduction in 3,3’,4,4’,5-pcnb is shown under the column ”Effect”, inference is given in the standard error (SE) and confidence interval columns. ”Modifier” is the region in the covariate space and its complement and Fold is the fold number.
Table 1: Effect Modification K-Fold Results
Exposure Effect SE Lower CI Upper CI Modifier Fold
LBXPCBLA 0.7047676 0.04008078 0.6262 0.7833 age_cent $\leq$ 14 1
LBXPCBLA 0.5625930 0.08915217 0.3879 0.7373 age_cent $>$ 14 1
LBXPCBLA 0.8017565 0.03205413 0.7389 0.8646 age_cent $\leq$ 11 2
LBXPCBLA 0.7246825 0.08895869 0.5503 0.8990 age_cent $>$ 11 2
LBXPCBLA 0.3222981 0.02627014 0.2708 0.3738 age_cent $\leq$ 11 3
LBXPCBLA 0.3271666 0.07242882 0.1852 0.4691 age_cent $>$ 11 3
LBXPCBLA -0.6316206 0.02034907 -0.6715 -0.5917 age_cent $\leq$ 11 4
LBXPCBLA -0.3153549 0.03551484 -0.3850 -0.2457 age_cent $>$ 11 4
LBXPCBLA 0.8372697 0.02145195 0.7952 0.8793 age_cent $\leq$ 14 5
LBXPCBLA 0.6842051 0.07540448 0.5364 0.8320 age_cent $>$ 14 5
LBXPCBLA 0.7044444 0.02454143 0.6563 0.7525 age_cent $\leq$ 11 6
LBXPCBLA 0.6348496 0.06230732 0.5127 0.7570 age_cent $>$ 11 6
LBXPCBLA 0.7364454 0.02360700 0.6902 0.7827 age_cent $\leq$ 17 7
LBXPCBLA 0.6057689 0.08581898 0.4376 0.7740 age_cent $>$ 17 7
LBXPCBLA 0.6412914 0.02309030 0.5960 0.6865 age_cent $\leq$ 11 8
LBXPCBLA 0.5364844 0.06194912 0.4151 0.6579 age_cent $>$ 11 8
LBXPCBLA 0.7782908 0.01962983 0.7398 0.8168 age_cent $\leq$ 11 9
LBXPCBLA 0.5685745 0.07544209 0.4207 0.7164 age_cent $>$ 11 9
LBXPCBLA 0.2118344 0.01881629 0.1750 0.2487 age_cent $\leq$ 17 10
LBXPCBLA 0.3519853 0.08829586 0.1789 0.5250 age_cent $>$ 17 10
7.7 Consistency of Results
Age consistently emerged as the maximum effect modifier across all folds. The exposure 3,3’,4,4’,5-pcnb (LBXPCBLA) was identified in 100% of the folds. Our findings indicate that a reduction in exposure to 3,3’,4,4’,5-pcnb by one standard deviation (52.7907) leads to an increase in LTL. This effect is more pronounced in younger populations, with central ages of 11, 14, and 17 consistently identified as split points. All these results were statistically significant based on the confidence intervals, underscoring the differential impact of this exposure across age.
7.8 Pooled Oracle Estimates
For each region and it’s complement, we provide the pooled oracle estimates. Here, $v$ denotes the regions for lower age and $v^{c}$ for higher age. Below is Table 2 of these pooled results.
Table 2: Effect Modification Region Pooled Results
Condition Psi Variance SE Lower CI Upper CI P-value
$v$ 0.4879 0.0001027 0.0101 0.5078 0.4680 0.0000
$v^{c}$ 0.3156 0.0010592 0.0325 0.2518 0.3794 3.11e-22
This indicates that the average impacts of shifting 3,3’,4,4’,5-pcnb by 1 standard deviation, -52.8, leads to an increase in telomere length of 0.488 in the lower age group, demarcated on average as centered age of 14, and of 0.316 in the higher age group.
We can compare the two groups based on $\Psi$ and the standard error using a two-sample z-test. For these results, the z-score is 5.06 resulting in a corresponding p-value p $<0.001$, indicating a significant difference between the two groups.
8 Conclusion
Our study presents a novel methodology, implemented in the open-source package, EffectXshift, found on github, to identify subpopulations that are differentially affected by chemical exposures and to estimate the effects of interventions within these subpopulations. Methodologically, our approach addresses several limitations of traditional causal inference methods by providing a data-driven, assumption-lean framework that leverages machine learning for both the estimation of exposure effects and the identification of effect modifiers. The use of stochastic shift interventions within our framework allows for more realistic and flexible modeling of exposure distributions, which is particularly relevant in environmental health studies, where exposures are often not binary or fixed.
By first establishing an oracle target parameter of interest as a definition of effect modification and then employing a data-adaptive target parameter strategy, we are able to deliver interpretable results grounded in theory. This approach is fundamentally different from other metalearning approaches in estimating heterogeneous treatment effects.
Applying this approach to the NHANES dataset, we demonstrated its ability to uncover significant effect modifiers, specifically highlighting age as a consistent modifier for 3,3’,4,4’,5-pcnb exposure and its association with leukocyte telomere length.
The findings of our analysis may have important implications for environmental health research and policy making. Identifying age as a significant effect modifier, our results suggest that younger individuals are more susceptible to the adverse effects of 3,3’,4,4’,5-pcnb exposure on telomere length. This information can inform targeted public health interventions aimed at reducing exposure to persistent organic pollutants (POPs) among vulnerable subpopulations, particularly younger individuals.
Despite the strengths of our approach, there are some limitations. The accuracy of our estimates depends on correct estimation of the clever covariate (ratio of densities), which can be difficult and the reason we offer an alternative approach via the classification approach. However, more research is necessary to ensure this reparameterization of the density estimation is sound in practice, although we show very little difference in results; our DGP is relatively simple. Additionally, while our method can handle complex and high-dimensional data, the computational cost may be significant for very large datasets. Future research should focus on further optimizing the computational efficiency of the EffectXshift package and exploring its applicability to other types of environmental exposures and outcomes.
In conclusion, the EffectXshift package provides a powerful tool for environmental health researchers to identify and estimate the differential impacts of chemical exposures between subpopulations. By facilitating the discovery of vulnerable groups and informing targeted interventions, this methodology contributes to the broader goal of reducing health disparities and improving public health outcomes.